# COMPASS: Cross-embOdiment Mobility Policy via ResiduAl RL and Skill Synthesis

### v1

@@ Abstract
As robots are increasingly deployed in diverse application domains, generalizable cross-embodiment mobility policies are increasingly essential. While classical mobility stacks have proven effective on specific robot platforms, they pose significant challenges when scaling to new embodiments. Learning-based methods, such as imitation learning (IL) and reinforcement learning (RL), offer alternative solutions but suffer from covariate shift, sparse sampling in large environments, and embodiment-specific constraints.
This paper introduces COMPASS, a novel workflow for developing cross-embodiment mobility policies by integrating IL, residual RL, and policy distillation. We begin with IL on a mobile robot, leveraging easily accessible teacher policies to train a foundational model that combines a world model with a mobility policy. Building on this base, we employ residual RL to fine-tune embodiment-specific policies, exploiting pre-trained representations to improve sampling efficiency in handling various physical constraints and sensor modalities. Finally, policy distillation merges these embodiment-specialist policies into a single robust cross-embodiment policy.
We empirically demonstrate that COMPASS scales effectively across diverse robot platforms while maintaining adaptability to various environment configurations, achieving a generalist policy with a success rate approximately 5X higher than the pre-trained IL policy. The resulting framework offers an efficient, scalable solution for cross-embodiment mobility, enabling robots with different designs to navigate safely and efficiently in complex scenarios.
@@
Figure 1: High-level overview of the COMPASS workflow: (1) Imitation learning produces a base policy and world model using readily available teacher policies on a mobile robot. (2) Residual RL fine-tunes the base policy for multiple embodiments, optimizing for physical constraints and sensor modalities. (3) Policy distillation consolidates these embodiment-specialist policies into one robust cross-embodiment policy. The screenshots illustrate successful deployments of the generalist policy across four different robot embodiments: Unitree G1 (humanoid), Unitree H1 (humanoid), Spot Mini (quadruped), and NOVA Carter (wheeled).
I Introduction
Robotics has experienced significant advances in both industry and daily life, driving the need for collaborative robots to handle increasingly complex tasks. However, developing robust cross-embodiment mobility policies remains challenging due to the many differences in morphological features, kinematics, and sensor configurations among diverse robotic platforms[1, 2, 3]. These discrepancies complicate efforts to create a universal policy that is both robust and adaptable in real-world settings.
@@ I Introduction
Classical mobility stacks [4, 5] excel on specific robots (especially wheeled platforms) but often require extensive retuning or redevelopment when ported to new embodiments with distinct sensor suites and physical constraints. This reliance on per-robot optimization drives interest in end-to-end learning approaches[6, 7], particularly for scaling across multiple robots.
Imitation learning (IL) is a compelling alternative, since it can leverage existing expert demonstrations and teacher policies. Despite its intuitive appeal, IL can succumb to covariate shift[8], where the policy encounters out-of-distribution states not seen during demonstrations. While advances in machine learning architectures[9, 10, 11] and data augmentation techniques[12] help mitigate these issues, adding more robot-specific factors increases data requirements and training complexity. Generating high-quality demonstrations for complex modalities (e.g., humanoids) further complicates pure IL approaches.
Reinforcement learning (RL) [13] offers another standard path to obtaining embodiment-specific policies, especially for tasks like locomotion. Yet navigational RL remains limited by large search spaces and sparse rewards in natural environments. Residual RL [14, 15] addresses these issues by refining a pre-trained policy in a data-driven manner, leading to faster convergence and greater stability. In parallel, emerging Visual-Language Action (VLA)[3, 1, 16, 17] models have shown promise for cross-platform tasks but typically operate through a low-dimensional waypoint-based action space or open-loop planning stages, making them less effective for platforms with higher-dimensional dynamics.
@@
To overcome these obstacles, we propose a unified workflow for cross-embodiment mobility (Figure 1). We start with an IL-based model for a mobile robot, using teacher policies to train a shared representation of mobility priors for environmental states and robot actions. As presented in the X-Mobility approach [6], this pre-trained model divides into two components: (i) a world model that encodes environment dynamics, and (ii) a policy that learns to produce expert-like actions. We then apply residual RL, starting from the IL policy and incrementally refining it into an embodiment-specific specialist. Thanks to the strong prior of IL, the residual RL step requires substantially fewer interactions to integrate morphological constraints and sensor streams effectively. At the final stage, we gather data from each specialist and apply policy distillation to merge them into a single cross-embodiment model, informed by an embodiment embedding vector. This distilled policy inherits the expertise of each specialist, enabling it to adapt seamlessly across various platforms. With this workflow, COMPASS achieves a 5X increase in task success rate and a 3X improvement in travel efficiency compared to the pre-trained IL policy, demonstrating its effectiveness and generalizability.
@@
To our knowledge, this three-stage framework is the first end-to-end pipeline for cross-embodiment mobility that systematically addresses morphological diversity, domain shift, and data efficiency. Residual RL mitigates covariate shift by aligning the training distribution with the deployment setting, while IL pre-training enables effective RL training even with sparse samplings. Our key contributions are as follows:
@@
  • A three-step workflow for scalable cross-embodiment mobility, integrating imitation learning, residual RL, and policy distillation.
  • A visual RL approach that leverages a world model to incorporate crucial temporal and spatial information, enabling efficient fine-tuning for different embodiments.
@@
  • Extensive benchmarks demonstrating that the proposed approach achieves robust generalization across diverse robot platforms while preserving the adaptability needed to succeed in varied environments.
II Method
@@ II-A Problem Statement
II-B Step 1: Imitation Learning for Mobility Priors
@@
II-B1 Latent State Modeling
@@
where $\mathbf{a}_{t}^{\star}$ are the expert (teacher) actions. More details can be referred to [6].
II-B2 Policy Learning in Latent Space
@@
$\mathbf{r}_{t}$ $=f_{\theta}(\mathbf{g}_{t}),$ (1)
$\mathbf{p}_{t}$ $=\Phi(\mathbf{s}_{t},\mathbf{r}_{t}),$ (2)
$\mathbf{a}_{t}$ $=\pi_{\theta}^{\text{IL}}\bigl{(}\mathbf{p}_{t}\bigr{)},$ (3)
@@
where $\ell(\cdot)$ could be a simple regression loss. This yields an IL-based mobility priors. The world model helps the policy generalize to out-of-distribution states by predicting future observations and latent transitions, thereby providing a robust encoded representation for decision making.
II-B3 X-Mobility
We leverages the X-Mobility [6] as our base policy, which integrates an autoregressive world model (see Fig. 2) with a velocity-prediction policy model. The learned latent state $\mathbf{s}_{t}$ encapsulates environment dynamics and constraints, while the policy head combines this state with route information to generate velocity commands. X-Mobility’s strong generalization performance indicates that its learned representation readily adapts to different embodiments.
II-C Step 2: Residual RL to Fine-Tune Specialists
@@
II-C1 Residual Policy Setup
Let $\mathbf{a}^{\text{base}}_{t}\;=\;\pi_{\theta}^{\text{base}}(\mathbf{p}_{t})$ be the action from the IL baseline. We introduce a residual policy $\pi_{\phi}^{\text{res}}$ that takes $\mathbf{p}_{t}$ and outputs $\mathbf{a}^{\text{res}}_{t}$. The final action is
$\mathbf{a}_{t}\;=\;\mathbf{a}^{\text{base}}_{t}+\mathbf{a}^{\text{res}}_{t}.$
@@ II-C1 Residual Policy Setup
II-C2 Reward Design
@@
  • Goal completion: Large positive reward on reaching the destination.
@@ II-C3 Training Loop
II-C4 Residual Policy Network Architecture
The residual policy network employs the same world model as the IL policy. We copy the weights from the IL action policy and reinitialize only the final output layer to learn the residual component. This strategy ensures stable training and focuses the residual learning on compensating for embodiment-specific performance gaps. For the critic network, we adopt a standard multi-layer perceptron (MLP) that takes the same policy state as input.
By building on a robust pre-trained base policy, the residual RL framework mitigates typical sparse sampling challenge, thereby enabling faster convergence to high-performance policies for each specific embodiment.
II-D Step 3: Policy Distillation to Combine Specialists
@@
II-D1 Data Collection from Specialists
@@ II-D3 Embodiment Embedding
II-D4 Distilled Policy Network Architecture
@@ III Experimental Setup
III-A Training
III-A1 IL Base Policy
@@ III-A2 Residual RL
III-A3 Policy Distillation
To distill the learned specialists into a single unified policy, we record 320 trajectories per embodiment using the same environment as residual RL training. Each trajectory spans 128 steps, yielding approximately 40k frames per embodiment. We then perform policy distillation using 4 Nvidia H100 GPUs by aligning each specialist’s output distribution.
III-B Benchmark
III-B1 Metrics
@@ III-B2 Scenarios
IV Results
In this section, we want to address three primary questions:
@@ IV Results
TABLE I: Benchmark Results
Embodiment Model Warehouse Single Rack Warehouse Multi Rack Office Combined
@@
IV-A Does residual RL produce high-performance embodiment specialists?
To investigate whether residual RL can improve performance relative to the IL baseline (X-Mobility), we compared the specialist policies for each embodiment against X-Mobility, and the results are summarized in Table I. In all evaluations, the specialist policies significantly outperform X-Mobility across both success rate (SR) and mobility efficiency (WTT). The SR is improved with a factor ranging from $5$x to $40$x, and the WTT is improved $3$x on average. This performance boost suggests that residual RL effectively leverages environment semantics and embodiment-specific physics. In particular, by exploring each robot’s physical limits during training, the specialist can produce commands that maximize efficiency while maintaining balance and stability. Moreover, the effectiveness of RL fine-tuning on environment semantic understanding can further bridges the Sim2Real gap when supported by a Real2Sim environment [20].
A notable observation is that while X-Mobility performs best on the Carter robot, its zero-shot performance on other embodiments regress significantly. This aligns with expectations and underscores the need for embodiment-specific policy adaptations to ensure robustness. Compared to the original X-Mobility paper, we observe a performance drop for Carter in warehouse environments within Isaac Lab, primarily due to simplified routes and variations in image rendering quality. This motivates future improvements, such as incorporating a hierarchical mobility stack for more sophisticated routing and exposing the model to diverse rendering conditions during latent state training.
IV-B Can policy distillation produce a robust generalist?
We next evaluate whether policy distillation can merge the specialists into a single generalist policy without sacrificing overall performance. As shown in Table I, the generalist achieves performance comparable to, and sometimes exceeding, that of the specialist policies. This is especially evident on the G1 robot, likely due to increased diversity in training scenarios and embodiments during distillation. These results confirm that our distillation approach is effective and that building a robust generalist from multiple specialists is both feasible and efficient.
@@ IV-B Can policy distillation produce a robust generalist?
IV-C How does COMPASS compare to training from scratch?
Finally, we explored whether we could train a policy from scratch using the same RL setup with the latent state but without leveraging the IL base actions. We observed that this RL-from-scratch approach struggled to converge even after 1,000 episodes as in Fig. 7, indicating significantly lower training efficiency. In contrast, our residual RL method exhibited notably faster convergence, underscoring how leveraging an IL baseline resolves the sparsity issues that commonly plague RL in mobility tasks.
V Ablation Studies
V-A Residual RL
In this section, we present ablation studies evaluating the various design decisions in our residual RL approach. We exclude the Carter robot from these experiments as it’s cross-embodiment effect is less pronounced.
V-A1 Curriculum on Goal Distance
In our RL training setup, we currently do not use a curriculum. To investigate whether such a strategy could improve training efficiency, we introduced a curriculum based on goal distance. Rather than enforcing a minimum goal distance from the outset, we gradually increased the probability of applying this constraint as training progressed. This incremental approach was intended to gradually raise task difficulty and enable smoother convergence.
However, the results in Table II indicate that the curriculum-based approach did not outperform the baseline without a curriculum. A likely explanation is that the IL base policy already provides a strong initialization, and reducing exposure to challenging scenarios with the curriculum may have limited the policy’s capacity to explore and succeed on more difficult tasks.
TABLE II: Ablation Studies on Residual RL.
Embodiment Model Warehouse Single Rack Warehouse Multi Rack Office Combined

### v3

@@ Abstract
As robots are increasingly deployed in diverse application domains, enabling robust mobility across different embodiments has become a critical challenge. Classical mobility stacks, though effective on specific platforms, require extensive per-robot tuning and do not scale easily to new embodiments. Learning-based approaches, such as imitation learning (IL), offer alternatives, but face significant limitations on the need for high-quality demonstrations for each embodiment.
To address these challenges, we introduce COMPASS, a unified framework that enables scalable cross-embodiment mobility using expert demonstrations from only a single embodiment. We first pre-train a mobility policy on a single robot using IL, combining a world model with a policy model. We then apply residual reinforcement learning (RL) to efficiently adapt this policy to diverse embodiments through corrective refinements. Finally, we distill specialist policies into a single generalist policy conditioned on an embodiment embedding vector. This design significantly reduces the burden of collecting data while enabling robust generalization across a wide range of robot designs. Our experiments demonstrate that COMPASS scales effectively across diverse robot platforms while maintaining adaptability to various environment configurations, achieving a generalist policy with a success rate approximately 5X higher than the pre-trained IL policy on unseen embodiments, and further demonstrates zero-shot sim-to-real transfer.
@@
Figure 1: High-level overview of the COMPASS workflow: (1) Imitation learning produces a base policy and world model using readily available teacher policies on a mobile robot. (2) Residual RL fine-tunes the base policy for multiple embodiments, optimizing for physical constraints and sensor modalities. (3) Policy distillation consolidates these embodiment-specialist policies into one robust cross-embodiment policy.
1 Introduction
While robotics has made significant strides in both industry and daily life, achieving robust mobility across diverse robot embodiments remains a fundamental open challenge. Morphological differences in hardware, kinematics and sensing configurations [1, 2, 3] create large variations in dynamics and perception, complicating efforts to build a universal and adaptable mobility policy for the real world.
@@ 1 Introduction
Classical mobility stacks [4, 5] excel in specific robots, especially wheeled platforms, but often require extensive manual retuning or redevelopment when ported to new embodiments with distinct sensor suites and physical constraints. This reliance on per-robot engineering has motivated interest in end-to-end learning approaches, particularly imitation learning (IL) [6, 7], which aim to scale mobility policies across multiple robots by leveraging data-driven training rather than manual redesign.
However, scaling IL to diverse embodiments introduces a major bottleneck: each new robot typically requires its own set of high-quality demonstrations. As morphological and sensor differences increase, so does the need for additional embodiment-specific data. For complex modalities such as humanoids, collecting demonstrations can be prohibitively expensive or impractical. In parallel, IL can also suffer from the distribution shift [8], leading to failures when the policy encounters out-of-distribution states during deployment. While advances in machine learning architectures [9, 10, 11] and data augmentation techniques [12] help mitigate the distribution shift within a given embodiment, extending these methods to accommodate greater morphological diversity greatly amplifies data requirements and training complexity, further complicating the scalability of pure IL approaches.
Nevertheless, we observe that many key aspects of mobility are inherently shared across robot embodiments, such as environmental understanding, obstacle avoidance, and goal reaching. This observation motivates our central insight: it should be possible to build a cross-embodiment mobility policy by leveraging expert demonstrations from only a single embodiment.
@@
Based on this principle, we propose COMPASS, a unified three-stage workflow that combines imitation learning, residual reinforcement learning, and policy distillation, as shown in Figure 1. We first pre-train a mobility policy on a single embodiment using IL, leveraging expert demonstrations to train a shared representation of mobility priors over environmental states and robot actions. Following the X-Mobility approach[6], this pre-trained model is decomposed into two components: (i) a world model that encodes environmental dynamics, and (ii) a policy that learns to produce actions. To efficiently adapt this policy to various embodiments, we integrate residual reinforcement learning (RL)[13, 14, 15] into our workflow. Rather than training policies from scratch, residual RL enables each embodiment-specific specialist to refine the pre-trained policy through small corrective adjustments, improving closed-loop performance with fewer environment interactions. Finally, we collect data from these specialists and apply policy distillation to merge their expertise into a single cross-embodiment model, conditioned on an embodiment embedding vector. This workflow amortizes the demonstration effort, enabling robust generalization across diverse platforms while requiring expert demonstrations from only a single embodiment. Experiments show that COMPASS achieves a 5X increase in task success rate and a 3X improvement in travel efficiency compared to the IL-only policy, demonstrating both its effectiveness and scalability.
@@
To the best of our knowledge, this three-stage framework is the first end-to-end pipeline for cross-embodiment mobility that systematically addresses morphological diversity, domain shift, and data efficiency. Our key contributions are as follows:
@@
  • We propose a unified three-stage workflow that combines imitation learning, residual reinforcement learning, and policy distillation to achieve scalable, data-efficient cross-embodiment mobility from demonstrations collected on a single embodiment.
  • We develop a visual RL approach that leverages a world model to efficiently refine policies across embodiments by incorporating temporal and spatial structure.
@@
  • We demonstrate extensive benchmarks across diverse robot platforms, showing that the proposed approach achieves robust generalization while preserving the adaptability needed to succeed in varied environments and zero-shot sim-to-real transfer.
2 Method
@@ 2.1 Problem Statement
2.2 Step 1: Imitation Learning for Mobility Priors
@@
2.2.1 Latent State Modeling
@@
where $\mathbf{a}_{t}^{\star}$ are the expert (teacher) actions.
2.2.2 Policy Learning in Latent Space
@@
$\mathbf{r}_{t}$ $=f_{\theta}(\mathbf{g}_{t}),$
$\mathbf{p}_{t}$ $=\Phi(\mathbf{s}_{t},\mathbf{r}_{t}),$
$\mathbf{a}_{t}$ $=\pi_{\theta}^{\text{IL}}\bigl(\mathbf{p}_{t}\bigr),$
@@
where $\ell(\cdot)$ could be a simple regression loss. This yields an IL-based mobility priors. The world model helps the policy generalize to out-of-distribution states by predicting future observations and latent transitions, thereby providing a robust encoded representation for decision-making.
2.2.3 X-Mobility
We leverage the X-Mobility [6] as our base policy, which integrates an autoregressive world model (see Fig. 2) with a velocity-prediction policy model. The learned latent state $\mathbf{s}_{t}$ encapsulates environment dynamics and constraints, while the policy head combines this state with route information to generate velocity commands. X-Mobility’s strong generalization performance indicates that its learned representation readily adapts to different embodiments.
2.3 Step 2: Residual RL to Fine-Tune Specialists
@@
2.3.1 Residual Policy Setup
Let $\mathbf{a}^{\text{base}}_{t}\;=\;\pi_{\theta}^{\text{base}}(\mathbf{p}_{t})$ be the action from the IL baseline. We introduce a residual policy $\pi_{\phi}^{\text{res}}$ that takes residual policy state $\hat{\mathbf{p}}_{t}$ and outputs $\mathbf{a}^{\text{res}}_{t}$. The final action is
$\mathbf{a}_{t}\;=\;\pi_{\theta}^{\text{base}}(\mathbf{p}_{t})+\pi_{\theta}^{\text{res}}(\hat{\mathbf{p}}_{t}).$
@@ 2.3.1 Residual Policy Setup
2.3.2 Residual Policy State
To enable effective residual policy learning, we introduce a residual policy state $\hat{\mathbf{p}}_{t}=\mathcal{G}(\mathbf{p}_{t})$, which augments the base policy state $\mathbf{p}_{t}$ with additional task-relevant information. This design allows the residual policy to complement the base policy by incorporating inputs that were not considered during the base policy training. For example, if the base policy omits goal heading information, we can inject its embedding into the residual policy state, allowing the final composite policy to account for the desired heading and achieve more stable goal-reaching behavior. This modular formulation offers flexibility to extend the capabilities of the base policy without retraining it from scratch.
2.3.3 Residual Policy Network Architecture
The residual policy network reuses the same policy model as in the base policy. We initialize it by copying the weights of the base action policy and introducing a linear projection layer to align the dimensionality between the base policy state and the residual policy state. The final output layer is reinitialized to exclusively learn the residual correction. This initialization strategy stabilizes training and encourages the residual component to specifically address embodiment-specific discrepancies. The critic network is implemented as a standard multi-layer perceptron (MLP), taking the same residual policy state as input.
2.3.4 Reward Design
@@
  • Goal completion: Large positive reward on reaching the destination with zero velocities.
@@ 2.3.5 Training Loop
By building on a robust pre-trained base policy, the residual RL framework mitigates the sparse sampling challenge. This enables faster convergence to high-performance policies for each embodiment.
2.4 Step 3: Policy Distillation to Combine Specialists
@@
2.4.1 Data Collection from Specialists
@@ 2.4.3 Embodiment Embedding
2.4.4 Distilled Policy Network Architecture
@@ 3 Experimental Setup
3.1 Training
3.1.1 IL Base Policy
@@ 3.1.2 Residual RL
3.1.3 Policy Distillation
To distill the learned specialists into a single unified policy, we record 320 trajectories per embodiment using the same environment as residual RL training. Each trajectory spans 128 steps, yielding approximately 40k frames per embodiment. We then perform policy distillation training on 4 Nvidia H100 GPUs by aligning each specialist’s output distribution.
3.2 Benchmark
3.2.1 Metrics
@@ 3.2.2 Scenarios
4 Results
This section addresses four primary questions:
@@ 4 Results
  • Can COMPASS policy trained in simulation be applied to real robot platforms?
4.1 Does residual RL produce high-performance embodiment specialists?
To investigate whether residual RL can improve performance relative to the IL baseline (X-Mobility), we compared the specialist policies for each embodiment against X-Mobility, and the results are summarized in Table 1. In all evaluations, the specialist policies significantly outperform X-Mobility across both success rate (SR) and mobility efficiency (WTT). The SR is improved with a factor ranging from 55x to 4040x, and the WTT is improved 33x on average. This performance boost suggests that residual RL effectively leverages environment semantics and embodiment-specific physics. In particular, by exploring each robot’s physical limits during training, the specialist can produce commands that maximize efficiency while maintaining balance and stability.
A notable observation is that while X-Mobility performs best on the Carter robot, its zero-shot performance on other embodiments degrade significantly. This aligns with expectations and underscores the need for embodiment-specific policy adaptations to ensure robustness. Compared to the original X-Mobility paper, we observe a performance drop for Carter in warehouse environments within Isaac Lab, primarily due to simplified routes and variations in image rendering quality. This motivates future improvements, such as incorporating a hierarchical mobility stack for more sophisticated routing and exposing the model to diverse rendering conditions during latent state training.
Table 1: Benchmark Results
Emb. Model Warehouse Single Warehouse Multi Office Combined
@@ 4.1 Does residual RL produce high-performance embodiment specialists?
4.2 Can policy distillation produce a robust generalist?
We next evaluate whether policy distillation can merge the specialists into a single generalist policy without sacrificing overall performance. As shown in Table 1, the generalist achieves performance comparable to, and sometimes exceeding, that of the specialist policies. This is especially evident on the G1 robot, likely due to increased diversity in training scenarios and embodiments during distillation. These results confirm that our distillation approach is effective and that building a robust generalist from multiple specialists is both feasible and efficient.
@@ 4.2 Can policy distillation produce a robust generalist?
4.3 How does COMPASS compare to training from scratch?
We also explored whether we could train a policy from scratch using the same RL setup with the latent state but without leveraging the IL base actions. We observed that this RL-from-scratch approach struggled to converge even after 1,000 episodes as in Fig. 7, indicating significantly lower training efficiency. In contrast, our residual RL method exhibited notably faster convergence, underscoring how leveraging an IL baseline resolves the sparsity issues that commonly plague RL in mobility tasks.
@@ 4.3 How does COMPASS compare to training from scratch?
4.4 Can COMPASS Policy trained in simulation be applied to real robot platforms?
To evaluate the Sim2Real transferability, we deployed it on two real-world robotic platforms powered by the NVIDIA Jetson Orin: Carter and Unitree G1 (Fig. 1). Leveraging the world model latent state representation, along with Isaac Lab’s high-quality visual rendering and accurate physics simulation, both robots achieved robust zero-shot transfer, successfully navigating cluttered environments without any additional fine-tuning. With localization enabled by cuVSLAM[18], the policy attained a success rate of approximately 80% across 20 trials with randomly placed obstacles, consistently avoiding collisions and reaching the designated targets. We anticipate that RL fine-tuning focused on semantic understanding of the environment, particularly when combined with a Real2Sim pipeline [19], can further narrow the Sim2Real gap.
Moreover, onboard inference with TensorRT requires only $\sim$30 ms per step as in Table 2, demonstrating that the policy is computationally efficient and well-suited for real-time deployment.
Table 2: Inference Latency and GPU Memory Usage
Model P50(ms) P95(ms) GPU Memory(MB)
Distilled Generalist 29.3389 29.3767 422
5 Ablation Studies
5.1 Residual RL
In this section, we present ablation studies evaluating the various design decisions in our residual RL approach. We exclude the Carter robot from these experiments as its cross-embodiment effect is less pronounced.
Table 3: Ablation Studies on Residual RL.
Emb. Model Warehouse Single Warehouse Multi Office Combined